when a gas filled in a closed container is heated through 1degree C its pressure is increased by
0.4% then initial temp of the gas was
1
Expert's answer
2016-12-28T06:11:06-0500
Initial temperature was T0 = (T2 - T1)*100 %/0.4 % = 1 K/0.004 = 250 K